OpenStudy (anonymous):

coefficient of x^4 in the expansion of (3x-1)^10

OpenStudy (campbell_st):

the binomal expansion is \[(a + b)^n = ^nC_{r}(a)^{n - r} \times (b)^r\] so to get a power of 4... r = 6 so you have \[^{10}C_{6} \times (3x)^{10-6}\times (-1)^6\] hope it helps

OpenStudy (anonymous):

\[Tr+1=nCr a ^{n-r}b^r~\in ~the~expansion~of~\left\{ \left( a+b \right)^n \right\}\] Here~n=10,a=3x,b=-1 \[Tr+1=10Cr \left( 3x \right)^{10-r}\left( -1 \right)^r=10Cr~ 3^{10-r}~x ^{10-r}\left( -1 \right)^r\] put 10-r=4,r=10-4=6 find T6+1=T7 and then coefficient of x^4

OpenStudy (anonymous):

\[10C6=10C(10-6)=10C4=\frac{ 10*9*8*7 }{4*3*2*1 }=210\] \[(-1)^6=1\] solve